A paper drinking cup in the shape of a right circular cone is constructed from 125 square centimeters of paper. If the height of the cone is 10 centimeters, find the radius correct to the hundredth of a centimeter.
Answer by Boreal(15235)

Area=pi*r^2+ pi*r sqrt (h^2+r^2)
125-pi*r^2=pi*r*sqrt(h^2+r^2)
divide by pi
39.79-r^2=r* sqrt(h^2+r^2)
square both sides
1583.14-79.58r^2+r^4=r^2*(100+r^2)=100r^2+r^4; r^4 cancel
1583.14-179.58r^2=0
1583.14=179.58r^2
divide by 179.58 and take the square root.
r=2.97 cm.